# biblio.bib

### v2

@@ 14 Conclusions and Future Work
Our approach is, in principle, applicable to any number of points. However, in practice, there seems to be a computational bottleneck in the construction of the Gröbner bases of the ideal of the critical configurations variety. The extension of this method to the case of seven points is the subject of ongoing work.
@@ Acknowledgements
We thank Pedro Raigorodsky and Carlos Beltrán for useful discussions during the completion of this work. We thank ClusterUY for the infrastructure for the numerical experiments. Matías Valdés acknowledges support from a PhD grant from Agencia Nacional de Investigación e Innovación (ANII) and Comisión Académica de Posgrado (CAP). Leandro Bentancur acknowledges support from a PhD grant from CAP. Marcelo Fiori and Mauricio Velasco were partially supported by ANII grant FCE-1-2023-1-176172. We thank the reviewers and the Editor for their constructive feedback, which contributed to an improved revised manuscript.
[…]
@@ Appendix A Solutions of minimal primes for $n=6$
In this appendix we describe the new solutions obtained with the procedure of Section 11.3, for the case of six points.
A.1 Factor $15x_{45}-$ 1 - Complex 1
@@
$X_{C1}=\begin{pmatrix}1&-1&-x_{15}&x_{15}&x_{15}&-x_{15}\\ &1&x_{15}&-x_{15}&-x_{15}&x_{15}\\ &&1&x_{45}&x_{45}&x_{25}\\ &&&1&x_{25}&x_{45}\\ &&&&1&x_{45}\\ &&&&&1\\ \end{pmatrix},\quad\begin{array}[]{l}x_{45}=\frac{1}{5},\\ x_{25}=-\frac{7}{5},\\ x_{15}=\pm\frac{i}{\sqrt{5}}\end{array}.$
A.2 Factor $1925x_{45}^{2}+28x_{45}+$ 19 - Complex 2
@@
$\left\{\begin{array}[]{l}25x_{45}^{2}+28x_{45}+19=0,\quad 8x_{13}^{2}-5x_{13}x_{45}+x_{13}-x_{45}-3=0,\\ 20x_{35}^{2}+10x_{35}x_{45}+10x_{35}-x_{45}-3=0,\\ A:=2x_{35}+\frac{3}{8}x_{45}+\frac{1}{8},\quad B:=-x_{13}+\frac{5}{8}x_{45}-\frac{1}{8},\\ C:=-x_{35}-\frac{1}{2}x_{45}-\frac{1}{2},\quad D:=-2x_{35}-\frac{5}{8}x_{45}-\frac{7}{8}.\end{array}\right..$
A.3 Factor $13x_{45}+$ 1 - Real 1
@@
$X_{N1}=\begin{pmatrix}1&x_{45}&x_{02}&x_{02}&x_{45}&x_{45}\\ &1&x_{02}&x_{02}&x_{45}&x_{45}\\ &&1&x_{23}&x_{02}&x_{02}\\ &&&1&x_{02}&x_{02}\\ &&&&1&x_{45}\\ &&&&&1\\ \end{pmatrix},\quad\begin{array}[]{l}x_{45}=-1/3,\\ x_{02}=0,\\ x_{23}=-1\end{array}.$
A.4 Factor $45x_{45}+$ 4 - Real 2
@@
$X_{N2}=\begin{pmatrix}1&x_{45}&x_{02}&x_{02}&x_{04}&x_{04}\\ &1&x_{02}&x_{02}&x_{04}&x_{04}\\ &&1&x_{02}&x_{02}&x_{02}\\ &&&1&x_{02}&x_{02}\\ &&&&1&x_{45}\\ &&&&&1\\ \end{pmatrix},\quad\begin{array}[]{l}x_{45}=-4/5,\\ x_{04}=1/10,\\ x_{02}=-1/5\end{array}.$
A.5 Factor $125x_{45}-$ 1 - Real 3
@@
$X_{N3}=\begin{pmatrix}1&x_{45}&x_{45}&x_{03}&x_{45}&x_{05}\\ &1&x_{12}&x_{03}&x_{12}&x_{45}\\ &&1&x_{03}&x_{12}&x_{45}\\ &&&1&x_{03}&x_{03}\\ &&&&1&x_{45}\\ &&&&&1\\ \end{pmatrix},\quad\begin{array}[]{l}x_{45}=1/25,\\ x_{05}=-23/25,\\ x_{12}=-11/25,\\ x_{03}=-1/5\end{array}.$
A.6 Factor $x_{45}$ Real 4
@@
$X_{N4}=\begin{pmatrix}1&x_{01}&x_{45}&x_{45}&x_{45}&x_{01}\\ &1&x_{45}&x_{45}&x_{45}&x_{01}\\ &&1&x_{01}&x_{01}&x_{45}\\ &&&1&x_{01}&x_{45}\\ &&&&1&x_{45}\\ &&&&&1\\ \end{pmatrix},\quad\begin{array}[]{l}x_{45}=0,\\ x_{01}=-1/2\end{array}.$

### v3

@@ 4.4 Classification of critical configurations
4.5 Seven Points
As we mentioned in Section 2, for seven points there is no known solution for the Fekete problem in $S^{2}$. To the best of our knowledge, no solution is known for seven points in $S^{3}$. The solution is already known for $S^{4}$ [17][Theorem 1.9] and $S^{5}$ [26].
We were unable to apply our method directly to the case of seven points, because the computation of the required Gröbner bases did not terminate using the time and resources at our disposal. However, we were able to compute the desired Gröbner bases if the polynomial system is simplified by adding one additional constraint, namely the presence of at least one dipole. Our main result in this section is that the only critical configuration in $S^{2}$ having at least one dipole is the 1:5:1. Our numerical experiments support the conjecture that this configuration is also the optimal configuration without the additional constraint. Our theorem reduces the problem of proving the conjecture, namely that 1:5:1 is the solution of the Fekete problem for seven points in $S^{2}$, to showing that some energy minimizer must necessarily have at least one dipole.
4.5.1 Numerical Experiments
We consider the Fekete problem as formulated in Equation (2), with objective function given by the logarithmic energy of $n$ points:
$E_{\log}:=-\log(E)=-\sum_{i=1}^{n}\sum_{j=i+1}^{n}\log\left(\|w_{i}-w_{j}\|^{2}\right).$
To estimate the local minima of the logarithmic energy in $(S^{d-1})^{n}$, we apply Riemannian Gradient Descent (RGD) on the manifold given by the product of the $n$ spheres. We use the package Pymanopt [30] to define the manifold and to calculate the projection to the tangent space and the retraction to the manifold. The step size in the tangent space is calculated with our own implementation of the Armijo rule, with parameters $\tau=0.5$, $r=1\times 10^{-4}$ and $\alpha_{0}=0.1$ [12][Alg. 4.2, p. 62]. The iterations are stopped when the norm of the projected gradient is less than $1\times 10^{-5}$. Algorithm 1 gives a pseudocode of the implemented algorithm, which can be found in the notebook gradient_descent_pymanopt. For each sphere $S^{d-1}$ we perform multiple tests, each with a different random initial configuration. We increase the number of tests together with the dimension of the sphere, to take into account that the search space is getting bigger.
Algorithm 1 Riemannian Gradient Descent on the product of $n$ spheres $S^{d-1}$ with Armijo rule.
Input: Initial configuration $w^{0}\in\left(S^{d-1}\right)^{n}$, gradTol, $r>0$, $\tau>0$, $\alpha_{0}>0$, maxIters. Retraction and Projection operators $R_{w}$ and $P_{w}$, respectively.
Output: configuration $w^{k}$ with $\|P_{w^{k}}\left(\nabla E_{\log}(w^{k})\right)\|\leq\text{gradTol}$.
[…]
3:while $k<\text{maxIters and }\|\nabla_{T}E_{\log}\|>\text{gradTol}$ do
4:
5:   while $E_{\log}\left(R_{w^{k}}\left(-\alpha\nabla_{T}E_{\log}\right)\right)>E_{\log}(w^{k})-r\alpha\|\nabla_{T}E_{\log}\|^{2}$ do
6:    $\alpha\leftarrow\tau\alpha$ $\triangleright$ Update stepsize with Armijo rule.
7:   end while
8:
9:   $w^{k+1}\leftarrow R_{w^{k}}\left(-\alpha\nabla_{T}E_{\log}\right)$ $\triangleright$ Update configuration using stepsize $\alpha$.
10:
11:   $\alpha\leftarrow\frac{\alpha}{\tau}$ $\triangleright$ Initial Armijo stepsize for next iteration.
12:
13:   $\nabla_{T}E_{\log}\leftarrow P_{w^{k+1}}\left(\nabla E_{\log}(w^{k+1})\right)$ $\triangleright$ Projected gradient for next iteration.
14:
15:   $k\leftarrow k+1$
16:end while
For $S^{2}$ we perform 1000 tests, and in all cases we obtain a configuration with energy value: $E_{\log}=-16.3649557$, when rounding up to seven digits. For each test, we also check that the estimated solution has a dipole. This, together with Theorem 2 and the termination criteria of the algorithm, implies that each estimated configuration is of the kind 1:5:1. These results suggest the following conjecture, which already appears in the bibliography, although without the details of the numerical experiments in which it is based [8][Conjecture 11.1], [13][Table 1.1].
Conjecture 1.
Configuration 1:5:1 is the unique kind of solution of the Fekete problem for $n=7$ points in $S^{2}$.
For $S^{3}$ we perform 5000 tests, and in all cases we obtain a configuration with the following energy, when rounding up to seven digits: $E_{\log}=-17.1587805$. For each test, we check that the estimated configuration has the following dot product matrix (or a permutation of it):
$WW^{T}=\begin{pmatrix}1&-1&0&0&0&0&0\\ -1&1&0&0&0&0&0\\ 0&0&1&-1&0&0&0\\ 0&0&-1&1&0&0&0\\ 0&0&0&0&1&-\frac{1}{2}&-\frac{1}{2}\\ 0&0&0&0&-\frac{1}{2}&1&-\frac{1}{2}\\ 0&0&0&0&-\frac{1}{2}&-\frac{1}{2}&1\\ \end{pmatrix}$.
This configuration has two dipoles, formed by the pairs $(w_{1},w_{2})$, $(w_{3},w_{4})$, and a regular simplex given by three points $w_{5}$, $w_{6}$ and $w_{7}$. Also, the dipoles are orthogonal to each other and to the regular simplex. Based on the Föpl notation, we denote this configuration as $4_{S^{1}}\times 3_{S^{1}}$, where we use $\times$ since the configuration lies in $S^{1}\times S^{1}$. These experimental results suggest the following conjecture.
Conjecture 2.
Configuration $4_{S^{1}}\times 3_{S^{1}}$ is the unique kind of solution of the Fekete problem for $n=7$ points in $S^{3}$.
For $S^{4}$ it is known that the Fekete problem has exactly two local minima, with one of them being a spurious local minima [16][Theorem 1.1]. In particular, the global minima is given by “two orthogonal simplexes, an equilateral triangle and a regular tetrahedron, inscribed in a great circle and a great 3-D hypersphere” [17][Theorem 1.9]. The spurious local minima is given by a dipole and a regular simplex of five points, orthogonal to each other. Our numerical experiments recover both local minima. For this we perform 10.000 tests, and we obtain configurations with two different energy values, as shown in Table 13 in the row of $S^{4}$. For each energy value, $E_{\log}^{(1)}$ and $E_{\log}^{(2)}$, we obtain configurations with respective Cartesian coordinate matrices $W_{1}$ and $W_{2}$, which correspond to the known local minima, as can be seen in the associated dot product matrices:
$W_{1}W_{1}^{T}=\begin{pmatrix}1&-\frac{1}{3}&-\frac{1}{3}&-\frac{1}{3}&0&0&1\\ -\frac{1}{3}&1&-\frac{1}{3}&-\frac{1}{3}&0&0&0\\ -\frac{1}{3}&-\frac{1}{3}&1&-\frac{1}{3}&0&0&0\\ -\frac{1}{3}&-\frac{1}{3}&-\frac{1}{3}&1&0&0&0\\ 0&0&0&0&1&-\frac{1}{2}&-\frac{1}{2}\\ 0&0&0&0&-\frac{1}{2}&1&-\frac{1}{2}\\ 0&0&0&0&-\frac{1}{2}&-\frac{1}{2}&1\\ \end{pmatrix},\ W_{2}W_{2}^{T}=\begin{pmatrix}1&-1&0&0&0&0&0\\ -1&1&0&0&0&0&0\\ 0&0&1&-\frac{1}{4}&-\frac{1}{4}&-\frac{1}{4}&-\frac{1}{4}\\ 0&0&-\frac{1}{4}&1&-\frac{1}{4}&-\frac{1}{4}&-\frac{1}{4}\\ 0&0&-\frac{1}{4}&-\frac{1}{4}&1&-\frac{1}{4}&-\frac{1}{4}\\ 0&0&-\frac{1}{4}&-\frac{1}{4}&-\frac{1}{4}&1&-\frac{1}{4}\\ 0&0&-\frac{1}{4}&-\frac{1}{4}&-\frac{1}{4}&-\frac{1}{4}&1\\ \end{pmatrix}$.
Table 13: RGD for $n=7$ points in $S^{d-1}\subset\mathbb{R}^{d}$. Iterations and final energy, with the number of times RGD converges to that value (in percentage).
# RGD Iterations Energy values
$S^{d-1}$ # RGD tests min. mean max. $E_{\log}^{(1)}$ $E_{\log}^{(2)}$
$S^{2}$ 1.000 20 2921 3073 -16.3649557 (100%) N/A
$S^{3}$ 5.000 50 72.6 145 -17.1587805 (100%) N/A
$S^{4}$ 10.000 60 98.0 253 -17.4985786 (84.2%) -17.4806735 (15.8%)
$S^{5}$ 50.000 18 24.1 58 -17.7932551 (100%) N/A
Finally, for seven points in $S^{k}$, with $k\geq 5$, it is known that the problem has a unique global minima, given by the 6-simplex [26]. We did 50.000 tests for $k=5$, and the algorithm always converged to a configuration with the energy of the simplex: $\simeq-17.7932551$. This suggests that there are no spurious local minima in this case.
4.5.2 Dipole implies conjectured minima in $S^{2}$
We now prove our main result of this section.
Theorem 2.
For the Fekete problem with $n=7$ points in $S^{2}$, the only critical configuration having a dipole is the 1:5:1.
To prove our result, we work with the system of critical configurations in Cartesian coordinates, as obtained in Section 3.1, and before taking dot products. This is useful as we want to work in $S^{2}$, while the dot product formulation gives critical configurations in any sphere dimension. Specifically, we use the following system of equations, with variables $w_{i}\in\mathbb{R}^{3}$ and $z_{ij}\in\mathbb{R}$:
$w_{i}^{T}w_{i}=1,\ (n-1)w_{i}-\sum_{j=1,j\neq i}^{n}\left(w_{i}-w_{j}\right)z_{ij}=\vec{0},\ \forall\ i;\qquad z_{ij}\left(1-w_{i}^{T}w_{j}\right)=1,\ \forall\ i<j.$ (10)
Table 14 shows the number of variables and equations of this system for some values of $n$. Note that this system has orthogonal symmetries that we should remove to apply our method.
Table 14: Number of variables and equations of the system given by Equations (10) in $\mathbb{R}^{3}$.
$n$ 3 4 5 6 7 8
variables $3n+\frac{n(n-1)}{2}$ 12 18 25 33 42 52
equations $4n+\frac{n(n-1)}{2}$ 15 22 30 39 49 60
To search for critical configurations having a dipole, we fix $w_{7}=(-1,0,0)$ and $w_{6}=(1,0,0)$ in the previous system, forming a dipole between $w_{6}$ and $w_{7}$. To remove orthogonal symmetries, we fix the last coordinate of another point at zero: $w_{5}=(x,y,0)$. This way of removing orthogonal symmetries is an alternative to what we did in Section 3.2 when taking dot products, with the difference that now we remain in $S^{2}$.
[…]
Figure 1: Seven points. One way of obtaining configuration 1:5:1 when $w_{6}$ and $w_{7}$ are fixed forming a dipole, and $w_{5}$ is restricted to $z=0$.
Configuration 1:5:1 is a particular solution of the resulting polynomial system. Moreover, when counting permutations, configuration 1:5:1 appears as a different solution exactly 48 times. This is because the point $w_{5}$ may be chosen in two different ways: $w_{5}=(0,1,0)$ or $w_{5}=(0,-1,0)$ (see Figure 1). In both cases, the points $w_{1}$ to $w_{4}$ may be freely permuted in $4!=24$ ways. This shows that the system has at least 48 different solutions (without counting multiplicities), all of them permutations of configuration 1:5:1.
On the other hand, we may count the exact number of solutions of the system. For this we first find a Gröbner basis with msolve, and then pass it to Macaulay2 to calculate the radical of the ideal. Using Theorem 1, we conclude that the number of solutions associated to this last ideal is 48. This is an upper bound on the number of different solutions. As this upper bound matches the lower bound, we conclude that the only kind of critical configuration with a dipole is configuration 1:5:1, as we wanted to prove. We note that the number of solutions of the ideal, before taking the radical, is $192=48\times 4$; which means that each of the 48 solutions has multiplicity four. Table 15 shows the execution time and RAM needed to calculate a Gröbner basis with grevlex monomial order and rational coefficients. The resources needed to calculate the radical of the ideal with M2, with a precomputed Gröbner basis as input, are negligible.
Table 15: Ideal with a dipole for $n=7$ in $S^{2}$, using Cartesian coordinates. Time and RAM of msolve v0.94 to calculate a Gröbner basis of the ideal with grevlex and 40 threads.
Time (s) Basis
# vars. # eqs. Total Lift to $\mathbb{Q}$ RAM # pols # mons file size
35 47 4542 806 27.2 GB 198 2096 27 KB
4.6 Comparison of Formulations in $S^{2}$
In this section we compare the performance of the formulation in Cartesian coordinates, given by Equations 10, with the one in terms of dot products, given by Equations (7), (8) and (9). For the cartesian formulation, we remove orthogonal symmetry by fixing $w_{n}=(-1,0,0)$, and $w_{n-1}=(x,y,0)$. We also need to avoid a dipole between these two points, since this would produce an $S^{1}$ symmetry with respect to the dipole axis. For this we introduce an auxiliary variable $t$, and the condition:
$t\left(w_{n-2}^{(1)}-1\right)=1;$
where $w_{n-2}^{(1)}$ denotes the first coordinate of $w_{n-2}$. This implies $w_{n-2}^{(1)}\neq 1$. On the other hand, for the dot product formulation, we add equations to limit the search of critical configurations to $S^{2}\subset\mathbb{R}^{3}$. Specifically, we add all $4\times 4$ minors of the dot product matrix $X$, which is equivalent to the condition: $\text{rank}(X)\leq 3$. Tables 16 and 17 show the performance of each formulation when calculating the leading monomials of a Gröbner basis with grevlex monomial order, using msolve v.0.94 in 40 threads.
Table 16: Cartesian formulation in $S^{2}$. Computation of leading monomials of a Gröbner basis.
msolve Basis Ideal
$n$ # vars # eqs max deg time (s) RAM # pols # mons $\dim$ $\deg$
4 15 22 3 0.12 3.85 MB 23 76 0 8
5 22 30 3 1.17 3.84 MB 247 6609 0 120
6 30 39 3 3110 14.4 GB 5681 3.268.865 0 2688
7 39 49 3 NA
Table 17: Dot product formulation in $S^{2}$. Computation of leading monomials of a Gröbner basis.
msolve Basis Ideal
$n$ # vars # eqs max deg time (s) RAM # pols # mons $\dim$ $\deg$
4 12 23 4 0.10 3.67 MB 15 57 0 4
5 20 60 4 0.66 3.66 MB 130 4178 0 37
6 30 276 4 564 3.23 GB 2037 1.288.061 0 717
7 42 1295 4 NA
As can be seen from the previous results, for up to six points, the dot product formulation uses less space and time resources than the Cartesian formulation, and the number of solutions is almost 3.8 times less than in the Cartesian formulation. Although time and space resources could change with a change in the monomial order, the difference in the number of solutions will remain; a matter that is relevant when analysing the different solutions in subsequent stages.
Also, we note that counting the different permutations of a given “kind” of solution is more difficult in the Cartesian formulation. In particular because the same “kind” of solution could appear in different permutation orbits.
@@ 5 Conclusions and Future Work
For seven points, although we are not able to apply our method to obtain all critical configurations, we can apply it to prove that if the Fekete problem in $S^{2}$ has a solution with a dipole, then it is given by configuration 1:5:1. Furthermore, our numerical experiments suggest that 1:5:1 is indeed the minimum energy configuration.
Our approach is, in principle, applicable to any number of points. However, in practice, there seems to be a computational bottleneck in the construction of the Gröbner bases of the ideal of the critical configurations variety. The extension of this method to characterize optimal configurations for the case of seven points is the subject of ongoing work.
@@ Acknowledgements
We thank Pedro Raigorodsky and Carlos Beltrán for useful discussions during the completion of this work. We thank ClusterUY for the infrastructure for the numerical experiments. Matías Valdés acknowledges support from a PhD grant from Agencia Nacional de Investigación e Innovación (ANII) and Comisión Académica de Posgrado (CAP). Leandro Bentancur acknowledges support from a PhD grant from CAP. Marcelo Fiori and Mauricio Velasco were partially supported by ANII grant FCE-1-2023-1-176172. We are grateful to the reviewers and handling Program Committee members of ISSAC 2025 Conference for their constructive feedback, which contributed to an improved revised manuscript. We also thank the reviewers and the Editor of the Journal of Symbolic Computation for their positive evaluation and helpful remarks.
@@ Appendix A Solutions of minimal primes for $n=6$
In this appendix we describe the new solutions obtained with the procedure of Section 4.2.3, for the case of six points.
A.1 Factor $5x_{45}-1$ - Complex 1
@@
$X_{C1}=\begin{pmatrix}1&-1&-x_{15}&x_{15}&x_{15}&-x_{15}\\ &1&x_{15}&-x_{15}&-x_{15}&x_{15}\\ &&1&x_{45}&x_{45}&x_{25}\\ &&&1&x_{25}&x_{45}\\ &&&&1&x_{45}\\ &&&&&1\\ \end{pmatrix},\qquad\begin{array}[]{l}x_{45}=\frac{1}{5},\\ x_{25}=-\frac{7}{5},\\ x_{15}=\pm\frac{i}{\sqrt{5}}\end{array}.$
A.2 Factor $25x_{45}^{2}+28x_{45}+19$ - Complex 2
@@
$\left\{\begin{array}[]{l}25x_{45}^{2}+28x_{45}+19=0,\qquad 8x_{13}^{2}-5x_{13}x_{45}+x_{13}-x_{45}-3=0,\\ 20x_{35}^{2}+10x_{35}x_{45}+10x_{35}-x_{45}-3=0,\\ A:=2x_{35}+\frac{3}{8}x_{45}+\frac{1}{8},\qquad B:=-x_{13}+\frac{5}{8}x_{45}-\frac{1}{8},\\ C:=-x_{35}-\frac{1}{2}x_{45}-\frac{1}{2},\qquad D:=-2x_{35}-\frac{5}{8}x_{45}-\frac{7}{8}.\end{array}\right.$.
A.3 Factor $3x_{45}+1$ - Real 1
@@
$X_{N1}=\begin{pmatrix}1&x_{45}&x_{02}&x_{02}&x_{45}&x_{45}\\ &1&x_{02}&x_{02}&x_{45}&x_{45}\\ &&1&x_{23}&x_{02}&x_{02}\\ &&&1&x_{02}&x_{02}\\ &&&&1&x_{45}\\ &&&&&1\\ \end{pmatrix},\qquad\begin{array}[]{l}x_{45}=-1/3,\\ x_{02}=0,\\ x_{23}=-1\end{array}.$
A.4 Factor $5x_{45}+4$ - Real 2
@@
$X_{N2}=\begin{pmatrix}1&x_{45}&x_{02}&x_{02}&x_{04}&x_{04}\\ &1&x_{02}&x_{02}&x_{04}&x_{04}\\ &&1&x_{02}&x_{02}&x_{02}\\ &&&1&x_{02}&x_{02}\\ &&&&1&x_{45}\\ &&&&&1\\ \end{pmatrix},\qquad\begin{array}[]{l}x_{45}=-4/5,\\ x_{04}=1/10,\\ x_{02}=-1/5\end{array}.$
A.5 Factor $25x_{45}-1$ - Real 3
@@
$X_{N3}=\begin{pmatrix}1&x_{45}&x_{45}&x_{03}&x_{45}&x_{05}\\ &1&x_{12}&x_{03}&x_{12}&x_{45}\\ &&1&x_{03}&x_{12}&x_{45}\\ &&&1&x_{03}&x_{03}\\ &&&&1&x_{45}\\ &&&&&1\\ \end{pmatrix},\qquad\begin{array}[]{l}x_{45}=1/25,\\ x_{05}=-23/25,\\ x_{12}=-11/25,\\ x_{03}=-1/5\end{array}.$
A.6 Factor $x_{45}$ - Real 4
@@
$X_{N4}=\begin{pmatrix}1&x_{01}&x_{45}&x_{45}&x_{45}&x_{01}\\ &1&x_{45}&x_{45}&x_{45}&x_{01}\\ &&1&x_{01}&x_{01}&x_{45}\\ &&&1&x_{01}&x_{45}\\ &&&&1&x_{45}\\ &&&&&1\\ \end{pmatrix},\qquad\begin{array}[]{l}x_{45}=0,\\ x_{01}=-1/2\end{array}.$